# PFML: Self-Supervised Learning of Time-Series Data Without Representation Collapse

### v2

@@ 4.1 PFML, MAE, and data2vec Experiments
For PFML pre-training, our models consist of a modality-specific frame-level encoder (detailed in Sections 4.2, 4.3, and 4.4 for IMU, speech, and EEG data, respectively) and a Transformer network consisting of $T$ Transformer encoder blocks. Between the encoder and Transformer networks there is a CNN-based relative positional encoder followed by a GeLU (Hendrycks & Gimpel, 2016) activation and layer normalization (Ba et al., 2016). We frame our input signals before feeding the data into an encoder model, and we compute functionals from these frames as our training targets. For multi-channel data, we compute functionals separately for each channel. The functionals are then z-score normalized across the entire pre-training dataset. For computational efficiency, we pre-compute the functionals of each signal frame before the pre-training process. After the Transformer encoder blocks, we add a linear projection to convert the Transformer outputs into predicted functionals. After pre-training, this linear projection is discarded. Pre-training is run until validation loss convergence, and we use the model with the lowest validation loss as our pre-trained model. Starting from an initial learning rate, we gradually reduce the learning rate during model training with a reduction factor of 0.5 based on the plateauing of the validation loss.
@@
We pre-train our models using MAE and data2vec in a similar manner as for PFML, and we use the same model architecture for all three pre-training algorithms. MAE pre-training is run in a similar manner as PFML pre-training, with the only exception of predicting the input signal frames instead of functionals. For data2vec pre-training, we used the instance-normalized (Ulyanov et al., 2016) and averaged outputs of each feed-forward part of all Transformer encoder blocks as our training targets. If we observed that a representation collapse occurred during data2vec pre-training, we restarted the pre-training process. For further details on the data2vec algorithm, see Baevski et al. (2022). We used mean squared error loss for all pre-training processes except for PFML with speech data, where we found L1 loss to work better.
@@
We fine-tune our pre-trained models in two stages. In the first stage, two randomly initialized fully-connected GeLU layers followed by a softmax function are added after the Transformer model. Then, these layers are fine-tuned separately as the weights of the encoder and Transformer are frozen. In the second stage, the entire model is fine-tuned with the same hyperparameters as in the first fine-tuning stage with one exception: The learning rate $LR$ is linearly increased from $0.001\cdot LR$ to $LR$ during a warm-up period of 20 training epochs, followed by reduction by a factor of 0.5 based on validation loss plateauing. We use weighted categorical cross-entropy loss by weighting the loss of each sample by its class’ inverse frequency.
@@
We also test the linear separability of the features learned by our pre-trained models. In this case, we only add one linear layer after the Transformer model, and we fine-tune this single layer while the weights of the encoder and Transformer are frozen. As a baseline, we perform the same linear evaluation for a randomly initialized model without any pre-training.
For pre-training, we use the RAdam (Liu et al., 2020), AdamW (Loshchilov & Hutter, 2019), and Adam (Kingma & Ba, 2015) optimizers for PFML, MAE, and data2vec, respectively. For fine-tuning, we use the Adam optimizer. For the “no pre-training” condition, we simply omit pre-training, the first fine-tuning stage, and the learning rate warm-up period of the second fine-tuning stage. We used an NVIDIA Tesla V100 GPU to train our models, and we implemented the code using PyTorch version 1.13.1. Our implementation is publicly available on GitHub. [^2] https://github.com/SPEECHCOG/PFML
@@
4.2 Infant Posture and Movement Classification
@@
For infant posture and movement classification, we use the multi-sensor IMU data from Airaksinen et al. (2022). The data contains 24-channel signals from infants (three gyroscope and three accelerometer channels, four limbs) with a sampling rate of 52 Hz. We window the signals into 120-sample frames (approx. 2.3 seconds) with 50% overlap. For further details about the dataset, see Airaksinen et al. (2022).
@@
For model pre-training, we use a 387-hour set of unlabeled IMU data from infant free-form play that has been automatically screened for signal quality (Vaaras et al., 2023b). This subset contains 4669 sequences of 260 consecutive frames, each corresponding to five minutes of data. As the encoder, we use the same four-layer CNN-based encoder architecture as in Airaksinen et al. (2022) with three minor modifications that were found to improve training efficiency and system performance when replicating the experiments of Airaksinen et al. (2022): We added layer normalization after the last two convolutions to make the pre-training process more stable, the kernel size of the second convolutional layer of the CNN encoder was changed from [4,5] to [3,5], and the originally temporally asymmetrical padding was set to [1,2] to make it symmetric. The pre-training data is randomly split into a training and validation set in a ratio of 80:20 sequences, and we input 260-frame sequences into the model.
@@
For fine-tuning our pre-trained models, we use a 29-hour (91,449 frames) labeled dataset of IMU data (41 recordings and distinct participants) for two separate tasks: posture classification and movement classification. The data contains nine annotated movement categories (still, roll left/right, pivot left/right, proto/elementary/fluent movement, transition) and seven annotated posture categories (prone, supine, left/right side, crawl posture, sitting, standing) for each 2.3-second frame. For model training, we use all annotated data, but we only use the frames in which all annotators agreed on the label for model testing. We train our models separately for both classification tasks using the so-called iterative annotation refinement labels from Airaksinen et al. (2020).
@@
Model fine-tuning is run using recording-level 10-fold cross-validation on the 41 distinct recordings of the labeled dataset. We split each training fold into separate training and validation sets in a ratio of 80:20 recordings. The unweighted average F1 score (UAF1) on the validation set is used as the training criterion, and we select the best-performing model based on validation set UAF1 score. We use random sensor dropout ($p=0.3$) for data augmentation during model fine-tuning. The final UAF1 score of fine-tuning is computed from an aggregate confusion matrix across all test folds. For further details regarding the pre-training and fine-tuning hyperparameters, see Appendix B.
4.3 Speech Emotion Recognition
@@
We use the 56-hour subset of Finnish speech of the NICU-A corpus (Vaaras et al., 2023a) for our speech emotion recognition experiments. This subset contains 129,007 utterances with a sampling rate of 16 kHz, of which 5198 and 345 belong to annotated training and testing sets, respectively. Each annotated utterance in NICU-A contains binary labels for emotional valence (positive/non-positive) and arousal (high/low). We window each speech signal into 30-ms frames with a 20-ms overlap. Each sequence is z-score normalized, and we zero-pad or truncate each normalized sequence into 3-second segments (301 frames). See Vaaras et al. (2023a) for further details on NICU-A.
@@
For model pre-training, we use all 129,007 utterances, and we input 301-frame sequences to our model. We use a four-layer CNN encoder with output channels $[128,128,128,128]$, kernel sizes $[10,8,4,4]$, strides of $[5,4,2,2]$, and paddings of $[3,2,1,1]$. Each layer is followed by layer normalization, a GeLU nonlinearity, and dropout. The last CNN layer is followed by average pooling with a kernel size of 6 before dropout. The pre-training utterances are randomly split into a training and validation set in a ratio of 80:20 sequences.
@@
We fine-tune and test our models separately for both classification tasks (valence/arousal) using the labeled 5198- and 345-utterance training and testing sets, respectively. The training set is randomly split into a training and validation set in a ratio of 80:20 utterances, and we select the best-performing model of the fine-tuning process based on the unweighted average recall (UAR) performance score on the validation set. This model is then used to compute the UAR performance score of the test set. See Appendix B for further details regarding the pre-training and fine-tuning hyperparameters.
4.4 Sleep Stage Classification
@@
For sleep stage classification, we use the pre-processed expanded Sleep-EDF Database (Kemp et al., 2000; Goldberger et al., 2000) from a study by Eldele et al. (2021). The dataset contains 30-second segments of the Fpz-Cz channel with a sampling rate of 100 Hz, comprising a total of 195,479 segments of EEG data. Each 30-second segment belongs to one of five annotated categories: wake, rapid eye movement (REM), non-REM stage 1, non-REM stage 2, or non-REM stages 3 and 4 combined. We z-score normalize each 30-second segment, and we window each segment into 4-second frames with 2 seconds of overlap, resulting into 14 frames for each segment.
@@
We fine-tune our models for sleep stage classification using 10-fold cross-validation at the test subject-level on the 78 test subjects of the dataset. Each training fold is split into training and validation sets at the test subject-level in a ratio of 80:20 test subjects. Similar to Sec. 4.2, we use the validation UAF1 score as our training criterion, and the testing UAF1 score is computed from an aggregate confusion matrix across all test folds. For further details on the training hyperparameters, see Appendix B.
4.5 TS2Vec Experiments
We also conducted a similar set of experiments as described in Section 4.1 for the TS2Vec method. TS2Vec stands out from the other three SSL methods used in the present experiments due to its hierarchical contrastive learning algorithm, which is integrated into the pre-defined CNN encoder architecture. Therefore, TS2Vec does not allow the use of the same modality-specific encoders used with PFML, MAE, and data2vec in our experiments. Additionally, it is not possible to pre-train the classification model (i.e. Transformer) in TS2Vec. Consequently, for all three data modalities (IMU, speech, and EEG data), we use the pre-defined TS2Vec encoder, and we do not pre-train the Transformer network. In order to compare the linear separability of the SSL features, we add one linear layer directly after the TS2Vec pre-trained encoder, and we fine-tune this single linear layer while the weights of the encoder are frozen.
For TS2Vec pre-training, we used the original TS2Vec implementation [^3] https://github.com/zhihanyue/ts2vec with default hyperparameter settings, except for two modifications: First, we observed that the pre-defined number of minibatch updates was insufficient for our pre-training experiments. We found it more effective to run TS2Vec pre-training similarly to PFML, MAE, and data2vec, i.e. until validation loss convergence, and then use the model with the lowest validation loss as the final pre-trained model. Second, we adjusted the output dimensionality of the TS2Vec encoder to match the modality-specific input dimensionalities of the Transformer (see Table 5 of Appendix B).
For fine-tuning the TS2Vec pre-trained models, we added a randomly initialized Transformer model and two randomly initialized fully-connected GeLU layers followed by a softmax function after the TS2Vec pre-trained encoder. The models were then fine-tuned similarly to the process described in Section 4.1, except that in the first fine-tuning stage, both the Transformer and the GeLU layers were fine-tuned while the weights of the encoder were frozen. The fine-tuning hyperparameters were the same as those used for the other pre-training algorithms (see Table 6 of Appendix B).
To demonstrate fair comparison, we also conducted three additional experiments using TS2Vec: First, similar to Section 4.1, we tested the “no pre-training” condition of the neural network architecture combining the TS2Vec encoder and Transformer classifier by fine-tuning the entire model at once using a randomly initialized model. Second, we tested the linear separability of a randomly initialized TS2Vec encoder by adding a single linear layer after the encoder and fine-tuning only this layer. Third, since the authors of the TS2Vec paper (Yue et al., 2022) used an SVM classifier with an RBF kernel for the SSL features in classification tasks, we adopted the same procedure for our experimental pipeline. All of the TS2Vec-related results are presented in Appendix F.
5 Results
Table 1 presents the fine-tuning results of the comparison of our PFML method against MAE, data2vec, TS2Vec, and not using pre-training at all. Across all three data modalities and five classification tasks, the end-use results show that PFML outperformed MAE and TS2Vec, and achieved results that are on par with data2vec. Using pre-training with any SSL method other than TS2Vec provided superior results as opposed to not using pre-training at all. Furthermore, for the classification of posture from IMU data, which is considered an easy task (Airaksinen et al., 2022), there were only minor differences in performance between all SSL methods other than TS2Vec. The comparably weak results for TS2Vec are most probably due to the fact that TS2Vec cannot utilize modality-specific encoders, and also since the Transformer model is not pre-trained during TS2Vec pre-training. Using SVM models instead of Transformers in TS2Vec classification, as in the original article of Yue et al. (2022), did not improve the results (see Appendix F). In sleep stage classification from EEG data, both MAE and PFML outperformed data2vec by a large margin. Also, the comparison between PFML and MAE showcases that it is more beneficial to predict functionals than to predict the input signal.
Table 2 shows the results of the linear evaluation experiments. Similar to the results of Table 1, PFML outperformed MAE and was comparable to data2vec when using the pre-trained models as feature extractors for linear classifiers. Again, both MAE and PFML outperformed data2vec by a large margin in sleep stage classification from EEG data. In valence classification from speech data and sleep stage classification from EEG data, the TS2Vec method excelled. However, for other classification tasks, the SSL features provided by TS2Vec were the weakest performance-wise. In the case of using a randomly initialized model as a feature extractor for linear classifiers, the classification accuracy was at chance-level in all cases except when classifying posture for IMU data.
@@ 5 Results
Table 1: Downstream task fine-tuning results for PFML, data2vec, MAE, TS2Vec, and not using pre-training at all for the five different classification tasks across the three different data modalities (IMU, speech, and EEG data).
44footnotetext: TS2Vec uses a different encoder compared to other experiments, since the TS2Vec algorithm is built into the pre-defined CNN encoder architecture.
@@
Table 2: Linear evaluation results for PFML, data2vec, MAE, TS2Vec, and a randomly initialized model.
55footnotetext: In TS2Vec, the linear layer is added directly after the pre-trained encoder.
@@
Table 3: Frequency of representation collapse across 10 runs of PFML, data2vec, MAE, and TS2Vec for each tested data modality.
The results of representation collapse experiments are shown in Table 3. As can be seen from the results, it is very common for representation collapse to occur with data2vec across all data modalities. On the contrary, the results indicate that PFML, MAE, and TS2Vec do not suffer from representation collapse: PFML and TS2Vec did not experience representation collapses at all, and MAE had a representation collapse only once. Furthermore, we attribute this single representation collapse of MAE to bad luck in model weight initialization, as in this particular case the model loss started diverging from the beginning of the pre-training process. The results showcase that methods like PFML, MAE, and TS2Vec, whose training targets inherently contain variance, are less prone to representation collapse compared to methods like data2vec that learn their own prediction targets.
6 Additional Hyperparameter Experiments
In order to demonstrate that it is more beneficial during pre-training to mask the latent features instead of masking the input directly, we ran PFML pre-training for all three datasets twice: either by masking the inputs or by masking the embeddings. Subsequently, we fine-tuned our models for all five classification tasks, and the results are shown in Table 7 of Appendix C. As can be observed from the results, it is more beneficial for downstream tasks if we alleviate the complexity of the pre-training task for the encoder by masking the embeddings instead of masking the inputs. The only exception was with EEG data, where it did not make a difference whether inputs or embeddings were masked.
For each data modality, we also experimented with different configurations of masking probability $p_{\text{m}}$ and the length of the masks $l_{\text{m}}$. We ran PFML pre-training using different configurations of $p_{\text{m}}$ and $l_{\text{m}}$, and then we fine-tuned the pre-trained models. For IMU and speech data, we only experimented with one classification task each, namely classification of movement from IMU data and classification of valence from speech data. The results for different configurations of $p_{\text{m}}$ and $l_{\text{m}}$ for IMU, speech, and EEG data are shown in Appendix C in Tables 8, 9, and 10, respectively. For IMU data, the differences between different masking strategies are rather small, whereas for speech and EEG data the selection of masking hyperparameters has a notable effect on fine-tuning performance.
We also experimented with the effect of discarding some of the functionals in PFML pre-training for IMU data. After pre-training, we fine-tuned our model for movement classification, and the results are presented in Table 11 of Appendix C. The results indicate that using the full set of 11 functionals during PFML pre-training provides the best outcome. As the number of discarded functionals increases, the prediction task becomes simpler and the training targets are able to capture less information of the input signal frames, leading to worse fine-tuning performance.
Finally, we tested different mask types for PFML pre-training using IMU data. We either replaced the masked embeddings with a fixed vector of zeros, ones, random Gaussian noise (as in e.g. Baevski et al. (2023)), or a learnable mask token (as in e.g. Baevski et al. (2022)). After PFML pre-training using the four different mask types, we fine-tuned the pre-trained models for movement classification. Table 12 of Appendix C presents the comparison results for different mask types. As can be observed, the choice between a mask of ones or random Gaussian noise does not have a notable impact on the performance. However, using a learnable mask token yielded slightly worse results than a vector of ones or random Gaussian noise, and a vector of zeros yielded the worst results. We observed that either using a vector of ones, random Gaussian noise, or learnable mask tokens for masking the embeddings promoted embedding variance, whereas using a vector of zeros provided a smaller level of variance for the embedding representations during pre-training. This lower level of variance for embeddings might potentially hinder the fine-tuning process, resulting into a lower performance in downstream tasks.
7 Conclusion
In this paper, we presented PFML, a novel SSL algorithm for time-series data that avoids the common SSL issue of representation collapse. PFML operates by predicting statistical functionals of the input signal corresponding to masked embeddings, given a sequence of unmasked embeddings. We demonstrated the effectiveness of PFML using five different classification tasks across three different data modalities: infant posture and movement classification from multi-sensor IMU data, emotion recognition from speech data, and sleep stage classification from EEG data. Our results show that PFML is superior to both a conceptually similar SSL method, MAE, and a contrastive learning-based SSL method, TS2Vec. Our results also show that PFML is on par with the current state-of-the-art data modality agnostic SSL method, data2vec, while being conceptually simpler and without suffering from representation collapse. The pros and cons for each of the SSL methods used in the present study are shown in Table 4. The fact that PFML matches the performance of data2vec while also avoiding the issue of representation collapse renders PFML more straightforward to apply to new time-series data domains, such as in the case of clinical time-series data. The present work may also be extended to other domains than time-series data, such as images where functionals could be computed of, e.g., image patches.

### v3

@@ IV-A PFML, MAE, and data2vec Experiments
For PFML pre-training, our models consist of a modality-specific frame-level encoder (detailed in Sections IV-B, IV-C, and IV-D for IMU, speech, and EEG data, respectively) and a Transformer network consisting of $T$ Transformer encoder blocks. Between the encoder and Transformer networks there is a CNN-based relative positional encoder followed by a Gaussian error linear unit (GELU) [47] activation and layer normalization [48]. We frame our input signals before feeding the data into an encoder model, and we compute functionals from these frames as our training targets. For multi-channel data, we compute functionals separately for each channel. The functionals are then z-score normalized across the entire pre-training dataset. For computational efficiency, we pre-compute the functionals of each signal frame before the pre-training process. After the Transformer encoder blocks, we add a linear projection to convert the Transformer outputs into predicted functionals. After pre-training, this linear projection is discarded. Pre-training is run until validation loss convergence, and we use the model with the lowest validation loss as our pre-trained model. Starting from an initial learning rate, we gradually reduce the learning rate during model training with a reduction factor of 0.5 based on the plateauing of the validation loss.
@@
We pre-train our models using MAE and data2vec in a similar manner as for PFML, and we use the same model architecture for all three pre-training algorithms. MAE pre-training is run in a similar manner as PFML pre-training, with the only exception of predicting the input signal frames instead of functionals. For data2vec pre-training, we used the instance-normalized [49] and averaged outputs of each feed-forward part of all Transformer encoder blocks as our training targets. If we observed that a representation collapse occurred during data2vec pre-training, we restarted the pre-training process. For further details on the data2vec algorithm, see Baevski et al. [8]. We used mean squared error loss for all pre-training processes except for PFML with speech data, where we found L1 loss to work better.
@@
We fine-tune our pre-trained models in two stages. In the first stage, two randomly initialized fully-connected GELU layers followed by a softmax function are added after the Transformer model. Then, these layers are fine-tuned separately as the weights of the encoder and Transformer are frozen. In the second stage, the entire model is fine-tuned with the same hyperparameters as in the first fine-tuning stage with one exception: The learning rate $\eta$ is linearly increased from $0.001\cdot\eta$ to $\eta$ during a warm-up period of 20 training epochs, followed by reduction by a factor of 0.5 based on validation loss plateauing. We use weighted categorical cross-entropy loss by weighting the loss of each sample by its class’ inverse frequency.
@@
We also test the linear separability of the features learned by our pre-trained models. In this case, we only add one linear layer followed by a softmax function after the Transformer model, and we fine-tune this single layer while the weights of the encoder and Transformer are frozen. As a baseline, we perform the same linear evaluation for a randomly initialized model without any pre-training.
@@
For pre-training, we use the RAdam [50], AdamW [51], and Adam [52] optimizers for PFML, MAE, and data2vec, respectively. For fine-tuning, we use the Adam optimizer. For the “no pre-training” condition, we simply omit pre-training, the first fine-tuning stage, and the learning rate warm-up period of the second fine-tuning stage. To demonstrate fair comparison, we carefully select the pre-training and fine-tuning hyperparameters for each SSL method separately in order to minimize the number of representation collapses during SSL pre-training, and to maximize the fine-tuning performance. For a complete list of pre-training and fine-tuning hyperparameters, refer to Appendix B in the supplementary material. We used an NVIDIA Tesla V100 GPU to train our models, and we implemented the code using PyTorch version 1.13.1. Our implementation is publicly available on GitHub. [^2] https://github.com/SPEECHCOG/PFML
IV-B Infant Posture and Movement Classification
@@
For infant posture and movement classification, we use the multi-sensor IMU data from Airaksinen et al. [53]. The data contains 24-channel signals from infants (three gyroscope and three accelerometer channels, four limbs) with a sampling rate of 52 Hz. We window the signals into 120-sample frames (approx. 2.3 seconds) with 50% overlap. For further details about the dataset, see Airaksinen et al. [53].
@@
For model pre-training, we use a 387-hour subset of a 1333-hour dataset of unlabeled IMU data [54]. This subset contains infant free-form play that has been automatically screened for signal quality, and it was selected for the present experiments based on the findings of Vaaras et al. [54], where using the subset yielded the best results in terms of SSL pre-training. This subset contains 4669 sequences of 260 consecutive frames, each corresponding to five minutes of data. As the encoder, we use the same four-layer CNN-based encoder architecture as in Airaksinen et al. [53] with three minor modifications that were found to improve training efficiency and system performance when replicating the experiments of Airaksinen et al. [53]: We added layer normalization after the last two convolutions to make the pre-training process more stable, the kernel size of the second convolutional layer of the CNN encoder was changed from $[4,5]$ to $[3,5]$, and the originally temporally asymmetrical padding was set to $[1,2]$ to make it symmetric. The pre-training data is randomly split into a training and validation set in a ratio of 80:20 sequences, and we input 260-frame sequences into the model.
@@
For fine-tuning our pre-trained models, we use a 29-hour (91,449 frames) labeled dataset of IMU data [53] (41 recordings and distinct participants) for two separate tasks: posture classification and movement classification. The data contains nine annotated movement categories (still, roll left/right, pivot left/right, proto/elementary/fluent movement, transition) and seven annotated posture categories (prone, supine, left/right side, crawl posture, sitting, standing) for each 2.3-second frame. For model training, we use all annotated data, but we only use the frames in which all annotators agreed on the label for model testing. We train our models separately for both classification tasks using the so-called iterative annotation refinement labels from Airaksinen et al. [55].
@@
Model fine-tuning is run using recording-level 10-fold cross-validation on the 41 distinct recordings of the labeled dataset. We split each training fold into separate training and validation sets in a ratio of 80:20 recordings. Due to large class imbalances in the labeled dataset of IMU data (see Airaksinen et al. [53]), the unweighted average F1 score (UAF1) is used as our performance metric. We use UAF1 on the validation set as the training criterion, and we select the best-performing model based on validation set UAF1 score. We use random sensor dropout ($p=0.3$) for data augmentation during model fine-tuning. The final UAF1 score of fine-tuning is computed from an aggregate confusion matrix across all test folds. For further details regarding the pre-training and fine-tuning hyperparameters, see Appendix B in the supplementary material.
IV-C Speech Emotion Recognition
@@
We use the 56-hour subset of Finnish speech of the NICU-A corpus [56] for our speech emotion recognition experiments. This subset contains 129,007 utterances with a sampling rate of 16 kHz, of which 5198 and 345 belong to annotated training and testing sets, respectively. Each annotated utterance in NICU-A contains binary labels for emotional valence (positive/non-positive) and arousal (high/low). We window each speech signal into 30-ms frames with a 20-ms overlap. Each sequence is z-score normalized, and we zero-pad or truncate each normalized sequence into 3-second segments (301 frames). See Vaaras et al. [56] for further details on NICU-A.
@@
For model pre-training, we use all 129,007 utterances, and we input 301-frame sequences to our model. We use a four-layer CNN encoder (slightly modified audio encoder of van den Oord et al. [4]) with output channels $[128,128,128,128]$, kernel sizes $[10,8,4,4]$, strides of $[5,4,2,2]$, and paddings of $[3,2,1,1]$. Each layer is followed by layer normalization, a GELU nonlinearity, and dropout. The last CNN layer is followed by average pooling with a kernel size of 6 before dropout. The pre-training utterances are randomly split into a training and validation set in a ratio of 80:20 sequences.
@@
We fine-tune and test our models separately for both classification tasks (valence/arousal) using the labeled 5198- and 345-utterance training and testing sets, respectively. The training set is randomly split into a training and validation set in a ratio of 80:20 utterances, and we select the best-performing model of the fine-tuning process based on the unweighted average recall (UAR) performance score on the validation set. This model is then used to compute the UAR performance score of the test set. See Appendix B in the supplementary material for further details regarding the pre-training and fine-tuning hyperparameters.
IV-D Sleep Stage Classification
@@
For sleep stage classification, we use the pre-processed expanded Sleep-EDF Database [57, 58] from a study by Eldele et al. [59]. The dataset contains 30-second segments of the Fpz-Cz channel with a sampling rate of 100 Hz, comprising a total of 195,479 segments of EEG data. Each 30-second segment belongs to one of five annotated categories: wake, rapid eye movement (REM), non-REM stage 1, non-REM stage 2, or non-REM stages 3 and 4 combined. We z-score normalize each 30-second segment, and we window each segment into 4-second frames with 2 seconds of overlap, resulting into 14 frames for each segment.
@@
We fine-tune our models for sleep stage classification using 10-fold cross-validation at the test subject-level on the 78 test subjects of the dataset. Each training fold is split into training and validation sets at the test subject-level in a ratio of 80:20 test subjects. Similar to Sec. IV-B, we use the validation UAF1 score as our training criterion, and the testing UAF1 score is computed from an aggregate confusion matrix across all test folds. For further details on the training hyperparameters, see Appendix B in the supplementary material.
@@
TABLE I: Downstream task fine-tuning results for PFML, data2vec, MAE, TS2Vec, and not using pre-training at all for the five different classification tasks across the three different data modalities (IMU, speech, and EEG data).
@@
TABLE II: Linear evaluation results for PFML, data2vec, MAE, TS2Vec, and a randomly initialized model.
@@
TABLE III: Frequency of representation collapse across 10 runs of PFML, data2vec, MAE, and TS2Vec for each tested data modality.
IV-E TS2Vec Experiments
We also conducted a similar set of experiments as described in Section IV-A for the TS2Vec method. TS2Vec stands out from the other three SSL methods used in the present experiments due to its hierarchical contrastive learning algorithm, which is integrated into the pre-defined CNN encoder architecture. Therefore, TS2Vec does not allow the use of the same modality-specific encoders used with PFML, MAE, and data2vec in our experiments. Additionally, it is not possible to pre-train the classification model (i.e. Transformer) in TS2Vec. Consequently, for all three data modalities (IMU, speech, and EEG data), we use the pre-defined TS2Vec encoder, and we do not pre-train the Transformer network. In order to compare the linear separability of the SSL features, we add one linear layer directly after the TS2Vec pre-trained encoder, and we fine-tune this single linear layer while the weights of the encoder are frozen.
For TS2Vec pre-training, we used the original TS2Vec implementation [^3] https://github.com/zhihanyue/ts2vec with default hyperparameter settings, except for two modifications: First, we observed that the pre-defined number of minibatch updates was insufficient for our pre-training experiments. We found it more effective to run TS2Vec pre-training similarly to PFML, MAE, and data2vec, i.e. until validation loss convergence, and then use the model with the lowest validation loss as the final pre-trained model. Second, we adjusted the output dimensionality of the TS2Vec encoder to match the modality-specific input dimensionalities of the Transformer (see Table 5 of Appendix B in the supplementary material).
For fine-tuning the TS2Vec pre-trained models, we added a randomly initialized Transformer model and two randomly initialized fully-connected GELU layers followed by a softmax function after the TS2Vec pre-trained encoder. The models were then fine-tuned similarly to the process described in Section IV-A, except that in the first fine-tuning stage, both the Transformer and the GELU layers were fine-tuned while the weights of the encoder were frozen. The fine-tuning hyperparameters were the same as those used for the other pre-training algorithms (see Table 6 of Appendix B in the supplementary material).
To demonstrate fair comparison, we also conducted three additional experiments using TS2Vec: First, similar to Section IV-A, we tested the “no pre-training” condition of the neural network architecture combining the TS2Vec encoder and Transformer classifier by fine-tuning the entire model at once using a randomly initialized model. Second, we tested the linear separability of a randomly initialized TS2Vec encoder by adding a single linear layer after the encoder and fine-tuning only this layer. Third, since the authors of the TS2Vec paper [35] used a support vector machine (SVM) classifier with a radial basis function kernel for the SSL features in classification tasks, we adopted the same procedure for our experimental pipeline. All of the TS2Vec-related results are presented in Appendix F in the supplementary material.
V Results
Table I presents the fine-tuning results of the comparison of our PFML method against MAE, data2vec, TS2Vec, and not using pre-training at all. Across all three data modalities and five classification tasks, the end-use results show that PFML outperformed MAE and TS2Vec, and achieved results that are on par with data2vec. Using pre-training with any SSL method other than TS2Vec provided superior results as opposed to not using pre-training at all. Furthermore, for the classification of posture from IMU data, which is considered an easy task [53], there were only minor differences in performance between all SSL methods other than TS2Vec. The comparably weak results for TS2Vec are most probably due to the fact that TS2Vec cannot utilize modality-specific encoders, and also since the Transformer model is not pre-trained during TS2Vec pre-training. Using SVM models instead of Transformers in TS2Vec classification, as in the original article of Yue et al. [35], did not improve the results (see Appendix F in the supplementary material). In sleep stage classification from EEG data, both MAE and PFML outperformed data2vec by a large margin. Also, the comparison between PFML and MAE showcases that it is more beneficial to predict functionals than to predict the input signal.
Table II shows the results of the linear evaluation experiments. Similar to the results of Table I, PFML outperformed MAE and was comparable to data2vec when using the pre-trained models as feature extractors for linear classifiers. Again, both MAE and PFML outperformed data2vec by a large margin in sleep stage classification from EEG data. In valence classification from speech data and sleep stage classification from EEG data, the TS2Vec method excelled. However, for other classification tasks, the SSL features provided by TS2Vec were the weakest performance-wise. In the case of using a randomly initialized model as a feature extractor for linear classifiers, the classification accuracy was at chance-level in all cases except when classifying posture for IMU data.